# Towards Practical Finite Sample Bounds for Motion Planning in TAMP

### v1

@@ 7 Conclusion
Our experiments show that despite being tighter than the original theoretical bound, our numerical sample bound is still too loose for practical use. In settings with a single narrow passage, we saw that the bound is tight within two orders of magnitude for low dimensions ($d=2$) and becomes increasingly loose as the number of dimensions grows. We then applied our numerical bound (for radius PRMs) as a heuristic to a TAMP planner equipped with a KNN PRM in problems where the narrow passage width was known. Our results show that only in low dimensional problems ($d=2$), the bound was a useful heuristic to help the planners scale the number of samples appropriately, given the complexity of the motion planning problem.
In the future, we wish to tighten our numerical bound for more practical use in higher dimensions. We can further explore Blumer et al.’s proof for improvements: our numerical algorithm only tightens Blumer’s bound over a single lossy step, and there are many more such steps that may lead to a significantly tighter bound. We also plan to finish extending the sample bound to KNN PRMs to bridge the gap between our theory and experiments, by investigating the connection between sampling to find $\alpha$-nets and effective connection radii.
@@ 7.0.1 Acknowledgements
We thank David Hsu, Sandeep Silwal, Justin Chen, Haike Xu, Josh Engels, Thomas Cohn, and Christopher Bradley for their invaluable support and feedback.
@@ A.1 PRM Pseudocode
Karaman and Frazzoli’s [15] implementation of the PRM algorithm uses two separate routines: a pre-processing routine that builds or grows a graph to a specified number of samples (Alg. 3), and a query routine that connects the starts and goals to the graph and runs a shortest-path search (Dijktra, etc., Alg. 4). For both algorithms, we assume several subroutines have already been provided:
  • SampleFree($\cdot$), which samples a collision-free point from the configuration space, and
Two implementations of the Near subroutine can be used: a range search that identifies the vertices within a radius of the query vertex or a $K$-nearest neighbor search.
@@ A.2 Proof of Lemmas 1 and 2 in Section 3
Let $L$ denote the length of $\sigma$. Let $\{B_{2\alpha}^{d}(p_{i})\}_{i\in[\left\lceil L/{2\alpha}\right\rceil]}$ denote a set of balls on $\sigma([0,1])$, where $B_{2\alpha,1}^{d}(p_{1})$ is centered on $x_{s}=p_{1}$, $B_{2\alpha}^{d}(p_{\left\lceil L/2\alpha\right\rceil})$ is centered on $x_{g}=p_{\left\lceil L/2\alpha\right\rceil}$. The remaining balls (for $i=2,\dots,\left\lceil L/{2\alpha}\right\rceil-1$) are centered along $\sigma([0,1])$, where the $p_{i}$ is spaced $2\alpha(i-1)$ away from $x_{s}$ along arclength of $\sigma$.
Next, we construct another set of balls, $\{B^{d}_{\alpha}(q_{j})\}_{j\in[\left\lceil L/2\alpha\right\rceil-1}]$, where $q_{j}$ is spaced $\alpha+2\alpha(j-1)$ along the arclength of $\sigma$ from $x_{s}=p_{1}$. We observe that there $q_{j}$ is spaced exactly between $p_{j}$ and $p_{j+1}$ along $\sigma$’s arclength.
Since these points are in a Euclidean space, we know that
$\left\lVert p_{j}-q_{j}\right\rVert,\left\lVert p_{j+1}-q_{j}\right\rVert\leq 2\alpha$
@@
Since $N$ forms an $\alpha$-net, we know that there must exist a point $x_{j}\in N\in B^{d}_{\alpha}(q_{j})$ for all $j$ ($x_{j}$ must be at most $\alpha$ from ball center $q_{j}$). Furthermore, we then know that the linear path between $x_{j}$ and $x_{j+1}$ is collision free for all $j$, since $x_{j},x_{j+1}\in B^{d}_{2\alpha}(p_{j+1})$ and balls are convex.
@@
We now describe the our proof for Lemma 2, Our arguments are centered around covering and point configurations and the probability they arise from random sampling processes, a topic well-studied at the intersection of statistics, combinatorics, and discrete geometry [2, 17, 22, 24] [^5] We will be using nomenclature from both computational geometry and statistics to avoid conflicting terminology..
Our intuition behind the definitions is grounded in trying to ‘hit’ a class of subsets of some space $X_{free}$ with a random finite set of points.
@@
$Y$ be a (potentially infinite) set, and let $\operatorname{\mathcal{F}}\subset 2^{Y}$. We call the tuple $(Y,\operatorname{\mathcal{F}})$ a range space. [^6] For those who are familiar with learning theory: another name for $\operatorname{\mathcal{F}}$ can be the hypothesis space of a set of binary classifiers or set of indicators associated with all elements of $\operatorname{\mathcal{F}}$.
For us, we will take the base space to be $X_{free}\subset\operatorname{\mathbb{R}}^{n}$, the obstacle-free configuration space of a robot. $\operatorname{\mathcal{F}}$ will be the set of $n$-dimensional balls that are a proper subset of $X_{free}$.
Often, hitting all the sets in $\operatorname{\mathcal{F}}$ with a finite number of points in $Y$ is too tall an order to ask. For our purposes, only hitting the ‘voluminous’ sets in $\operatorname{\mathcal{F}}$ is enough.
@@
Let $(Y,\operatorname{\mathcal{F}})$ is a range space. Suppose that $\mu$ is a measure on $Y$ such that every $S\in\mathcal{F}$ is measurable with respect to $\mu$.
@@
Intuitively, $\epsilon$ is a threshold on the minimum volume of $S\in\operatorname{\mathcal{F}}$ to be hit by a point in $N$. There could be many different $\epsilon$-transversals on $(Y,\operatorname{\mathcal{F}})$, and we prefer to find smaller $\epsilon$-transversals if we can.
Blumer et al. [2], generalizing Welzl and Haussler’s work [10] from discrete to continuous range spaces showed that random sampling can find $\epsilon$-transversals with high probability.
@@
$m\geq\max\left\{\frac{4}{\epsilon}\log\frac{2}{\gamma},\frac{8d}{\epsilon}\log\frac{13}{\epsilon}\right\},$
@@
Intuitively, the VC-dimension is a quantification of the complexity, or intricacy, of the ways the sets in $\operatorname{\mathcal{F}}$ intersect with $Y$ and each other. We now move to the technical definition:
@@
Let $(Y,\operatorname{\mathcal{F}})$ a range space. We say that a subset $A\subset Y$ is shattered by $\operatorname{\mathcal{F}}$ if each of the subsets of $A$ can be obtained as the intersection of some $S\in\operatorname{\mathcal{F}}$ with $A$. We define the VC-dimension of $\operatorname{\mathcal{F}}$ as the supremum of the sizes of all finite shattered subsets of $Y$. If arbitrary large subsets can be shattered, the VC-dimension is $\infty$.
Our arguments only require the fact that the VC-dimension of a set of $d$-dimensional (Euclidean) balls in $\operatorname{\mathbb{R}}^{d}$ is $d+1$ [22, 24]. When uniformly sampling from a space $X_{free}$, we are in a favorable scenario where $\epsilon$-transversal (hitting set of balls in $X_{free}$) $\epsilon$-net (covering $X_{free}$ with balls) defined in Def. 2 coincide.
@@
The backward direction is even shorter. Let $N$ be a $\epsilon^{d}\operatorname{\mathrm{P}}(B_{1}^{d})$-transversal. Let $x\in\mathcal{Y}$. Let $S$ be the $d$-dimensional ball centered at $x$. Then there must exist $x_{n}\in N$ such that $x_{n}\in S$, so $||x-x_{n}||\leq\epsilon$. $\blacksquare$
We now have all the ingredients we require for Lemma 2:
@@
By Lemma 4 $\epsilon^{d}\operatorname{\mathrm{P}}(B^{d}(1))$-transversals is an $\epsilon$-net. We apply Lemma A.1, and the result follows. $\blacksquare$
A.3 Proofs of Lemma 3 and Prop. 1 in Section 4
The proof of Lemma 3 is a simple proof of contradiction. The intuition for the proof is pictorially displayed in Fig. 3.
@@
We start with the first statement. Let $r\in\operatorname{\mathbb{R}}$ such that $0<r<\min_{v\in V}\left\lVert v-\operatorname{NN}_{k+1}(v)\right\rVert$. Suppose, for the sake of contradiction, that there exists $u,v\in V$ such that CollisionFree$(u,v)$ returns True but $(u,v)\notin E$.
Then $v$ must not be a $K$th nearest neighbor of $u$, or vice versa. Without loss of generality, assume the former, and so $v$ must be at least a $(K+1)$th nearest neighbor of $u$. But then $r<\min_{v\in V}\left\lVert v-\operatorname{NN}_{k+1}(v)\right\rVert<\left\lVert u-v\right\rVert$, so we have a contradiction.
The proof of the second statement is fast: suppose $r\geq\min_{v\in V}||v-\operatorname{NN}_{k+1}(v)||$. Let $v^{*}$ be the minimizing vertex in the expression above. Then $(v^{*},\operatorname{NN}_{k+1}(v))$ is a non-edge and $\left\lVert v^{*}-\operatorname{NN}_{k+1}(v^{*})\right\rVert\leq r$. $\blacksquare$
This proof is loosely inspired by the KNN proof for Lemma 58 done by Karaman and Fazzoli in [15], which uses a Chernoff bound to control the number of samples that hit a ball with a radius that scales with the volume and dimension of the environment and number of samples taken:
@@
We prove our bound by controlling the number of samples land within a ball of radius $r$ of another sample using a Chernoff bound.
@@
$\operatorname{\mathrm{P}}\left(\forall v\in V,\sum_{w\in V\backslash\{v\}}I_{v,w}(r)<K+1\right)$
@@
The second line takes a union bound over all vertices. The third line applies the bound $\operatorname{\mathrm{P}}(B_{r}^{d}(v)\cap X)\leq\operatorname{\mathrm{P}}(B_{r}^{d}(v))$. To obtain a tail estimate on the bimonial random variable above, we use the Chernoff bound stated in [3]:
@@
where $D_{KL}$ is the KL-divergence between two Bernoulli variables, and $p=\operatorname{\mathrm{P}}(B_{r}^{d})$ and $t=K/(n-1)-p$ (our bound on $r$ will avoid a degenerate $t\leq 0$).
Using the well-known inequality $D_{KL}(p_{1}||p_{2})\geq(p_{1}-p_{2})^{2}/2p$ when $p_{1}<\frac{1}{8}$ and $p_{1}>p_{2}$ (the mild technical condition stated in the hypothesis) we have that:
@@
What we have shown is that if we choose a ball radius $r$ as expressed in Equation 12, then each ball of radius $r$ centered at each vertex in the PRM contains less than $K+1$ other vertices with probability at least $1-\gamma$. By the construction of the event, we know that:
@@
A.4 Numerical Algorithm Results
@@ A.4.1 Analysis of Correctness of Algorithm 2
We begin by stating the following theorem that formalizes the algorithm’s correctness and sketching out the proof:
@@
We observe from the form of the expression (lower order polynomial and exponential on the left term, higher order polynomial and linear factor on the right term) and conclude two possibilities: (1) $f(m)<0$ for all $m\geq 1$ or (2) $f(m)\geq 0$ for some $k>m$, and then $f(m)<0$ for all $m>k$. If we are in the first case, we then write $k=0$ for ease of the rest of the analysis.
In either case, the doubling search of the numerical algorithm will double $N_{u}$ until $N_{u}>k$ and $f(N_{u})\leq\gamma$. Thus, we observe that the binary search will preserve two invariants: (1) $N_{l}\leq m^{*}\leq N_{u}$, since the search increases $N_{l}$ if either (1) $f(\left\lfloor(N_{l}+N_{u})/2\right\rfloor)>\gamma$ or (2) $N_{l}<k$ (by the decrease check on line 6), and (2) the intervals close by half each time.
Thus, since we are searching over a discrete set while the invariants hold, then the algorithm will return the correct answer.
A.4.2 Numerical Algorithm for Proposition 1 and Experimental Results

### v3

@@ 7 Conclusion
Our experiments show that despite being less conservative than the original theoretical bound, our numerical sample bound is still too loose for practical use. In settings with a single narrow passage using a radius PRM, we saw that the estimated number of samples by the numerical algorithm is tight within two to three orders of magnitude for planar problems. The estimate becomes increasingly loose and less useful as the number of dimensions grows. When the numerical algorithm is applied as a heuristic to estimate the number of samples of the KNN PRM, we found the algorithm to be more useful for low dimension problems ($d=2,3$) than as a guarantee for radius PRMs. We also applied the heuristic to guide sampling effort of a KNN PRM in a TAMP planner in problems where the narrow passage width was known. Our results show that only in planar problems, the bound was a useful heuristic to help the planners scale the number of samples appropriately to the sample complexity of the motion planning problem.
In the future, we wish to tighten our theoretical and numerical bounds for more practical use. We wish to introduce a version bound that has a more realistic characterization of the worst-case planning problem. At present, the ‘worst-case’ problem is a configuration space that consists of a single narrow passage with constant clearance along every dimension. In many problems, the clearances vary in the configuration space, with narrow widths in some but not all dimensions. Introducing additional parameters amenable to iterative search for a more realistic worst case may improve bound tightness for planning problems that appear in practice. We can further explore Blumer et al.’s proof for improvements: our numerical algorithm only tightens Blumer’s bound over a single lossy step, and there are many more such steps that may lead to a significantly tighter bound. We also plan to finish extending the sample bound to KNN PRMs to bridge the gap between our theory and experiments.
@@ 7.0.1 Acknowledgements
We thank David Hsu, Steven LaValle, Sandeep Silwal, Justin Chen, Haike Xu, Josh Engels, Thomas Cohn, Christopher Bradley, Michael Noseworthy, Erick Fuentes, and Chad Kessens for their invaluable support. This work was supported by the National Science Foundation Graduate Research Fellowship Program under Grant No. 2141064. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the National Science Foundation.
@@ A.1 PRM Pseudocode
Karaman and Frazzoli’s [16] implementation of the PRM algorithm uses two separate subroutines: a pre-processing subroutine that grows a graph to a specified number of samples (Alg. 3), and a query routine that connects the starts and goals to the graph and solves a shortest-path search (using Dijkstra’s algorithm, etc.; Alg. 4). For both algorithms, we assume several subroutines have already been provided:
  • SampleFree($\cdot$), which samples a collision-free point from the configuration space.
Standard implementations of the Near subroutine include a range search that identifies the vertices within a radius of the query vertex and a $K$-nearest neighbor search.
@@ A.2 Proof of Lemmas 1 and 2 in Section 3
Let $L$ denote the length of $\sigma$. Let $\{B_{2\alpha}^{d}(p_{i})\}_{i\in[\left\lceil L/{2\alpha}\right\rceil]}$ denote a set of balls on $\sigma([0,1])$, where $B_{2\alpha,1}^{d}(p_{1})$ is centered on $x_{s}=p_{1}$ and $B_{2\alpha}^{d}(p_{\left\lceil L/2\alpha\right\rceil})$ is centered on $x_{g}=p_{\left\lceil L/2\alpha\right\rceil}$. The remaining balls (for $i=2,\dots,\left\lceil L/{2\alpha}\right\rceil-1$) are centered along $\sigma([0,1])$, where the $p_{i}$ is spaced $2\alpha(i-1)$ away from $x_{s}$ along the arclength of $\sigma$.
Next, we construct another set of balls, $\{B^{d}_{\alpha}(q_{j})\}_{j\in[\left\lceil L/2\alpha\right\rceil-1}]$, where $q_{j}$ is placed $\alpha+2\alpha(j-1)$ along the arclength of $\sigma$ from $x_{s}=p_{1}$. We observe that $q_{j}$ is placed exactly between $p_{j}$ and $p_{j+1}$ along $\sigma$’s arclength. Since these points are in a Euclidean space, straight lines represent shortest paths, so
$\left\lVert p_{j}-q_{j}\right\rVert,\left\lVert p_{j+1}-q_{j}\right\rVert\leq 2\alpha,$
@@
Since $N$ forms an $\alpha$-net, we know that there must exist a point $x_{j}\in N\in B^{d}_{\alpha}(q_{j})$ for all $j$ ($x_{j}$ must be at most $\alpha$ from ball center $q_{j}$). Furthermore, we observe that the linear path between $x_{j}$ and $x_{j+1}$ is collision-free for all $j$, since $x_{j},x_{j+1}\in B^{d}_{2\alpha}(p_{j+1})$ and balls are convex.
@@
We now describe the our proof for Lemma 2. Our arguments are centered around covering point configurations and the probability they arise from random sampling processes, a topic well-studied at the intersection of statistics, combinatorics, and discrete geometry [^2] We will be using nomenclature from both computational/combinatorial geometry and statistics to avoid conflicting terminology. [2, 18, 24, 26]. Our intuition behind the definitions is grounded in trying to ‘hit’ a class of subsets of some space $X_{free}$ with a random finite set of points.
@@
$Y$ be a (potentially infinite) set and let $\operatorname{\mathcal{F}}\subset 2^{Y}$. We call the tuple $(Y,\operatorname{\mathcal{F}})$ a range space. [^3] For those who are familiar with learning theory: another name for $\operatorname{\mathcal{F}}$ can be the hypothesis space of a set of binary classifiers or set of indicators associated with all elements of $\operatorname{\mathcal{F}}$.
We will take the base space to be $X_{free}\subset\operatorname{\mathbb{R}}^{n}$, the obstacle-free configuration space of a robot. $\operatorname{\mathcal{F}}$ will be the set of $n$-dimensional balls that are a proper subset of $X_{free}$. Often, hitting all the sets in $\operatorname{\mathcal{F}}$ with a finite number of points in $Y$ is too tall an order to ask. Only hitting the ‘voluminous’ sets in $\operatorname{\mathcal{F}}$ is enough for our purposes.
@@
Let $(Y,\operatorname{\mathcal{F}})$ be a range space. Suppose that $\mu$ is a measure on $Y$ such that every $S\in\mathcal{F}$ is measurable with respect to $\mu$.
@@
Intuitively, $\epsilon$ is a threshold on the minimum volume of $S\in\operatorname{\mathcal{F}}$ to be hit by a point in $N$. There could be many different $\epsilon$-transversals on $(Y,\operatorname{\mathcal{F}})$, and we prefer to find smaller $\epsilon$-transversals if we can. Blumer et al. [2], generalizing Welzl and Haussler’s work [11] from discrete to continuous range spaces showed that random sampling can find $\epsilon$-transversals with high probability. We can derive the following statement by combining Theorem A2.1, Proposition A2.1, and Lemma A2.4 by Blumer et al. [2]:
@@
$m\geq\max\left\{\frac{4}{\epsilon}\log_{2}\frac{2}{\gamma},\frac{8d}{\epsilon}\log_{2}\frac{13}{\epsilon}\right\},$
@@
Intuitively, the VC-dimension quantifies the complexity, or intricacy, of how the sets in $\operatorname{\mathcal{F}}$ intersect with $Y$ and each other. We now move to the technical definition of the VC-dimension:
@@
Let $(Y,\operatorname{\mathcal{F}})$ be a range space. We say that a subset $A\subset Y$ is shattered by $\operatorname{\mathcal{F}}$ if each of the subsets of $A$ can be obtained as the intersection of some $S\in\operatorname{\mathcal{F}}$ with $A$. We define the VC-dimension of $\operatorname{\mathcal{F}}$ as the supremum of the sizes of all finite shattered subsets of $Y$. If arbitrarily large subsets can be shattered, the VC-dimension is $\infty$.
Our argument only requires that the VC-dimension of a set of $d$-dimensional (Euclidean) balls in $\operatorname{\mathbb{R}}^{d}$ is $d+1$ [24, 26]. When uniformly sampling from a space $X_{free}$, we are in a favorable scenario where the definitions of $\epsilon$-transversal (hitting set of balls in $X_{free}$) $\epsilon$-net (covering $X_{free}$ with balls) defined in Def. 2 coincide (but for different $\epsilon$’s).
@@
The backward direction is even shorter. Let $N$ be a $\epsilon^{d}\operatorname{\mathrm{P}}(B_{1}^{d})$-transversal. Let $x\in Y$. Let $S$ be the $d$-dimensional ball centered at $x$. Then there must exist $x_{n}\in N$ such that $x_{n}\in S$, so $||x-x_{n}||\leq\epsilon$. $\blacksquare$
We now have all the ingredients we require for Lemma 2.
@@
By Lemma 4, $\epsilon^{d}\operatorname{\mathrm{P}}(B^{d}(1))$-transversals is an $\epsilon$-net. We apply Lemma A.1, and the result follows. $\blacksquare$
A.3 Proofs of Lemma 3 and Prop. 1
The proof of Lemma 3 is a simple proof by contradiction. The intuition for the proof is pictorially displayed in Fig. 3.
@@
We start with the forward direction. Let $r\in\operatorname{\mathbb{R}}$ such that $0<r<\min_{v\in V}\left\lVert v-\operatorname{NN}_{k+1}(v)\right\rVert$. Suppose, for the sake of contradiction, that there exists $u,v\in V$ such that CollisionFree$(u,v)$ returns True but $(u,v)\notin E$. Then $v$ must not be a $K$th nearest neighbor of $u$, or vice versa. Without loss of generality, assume the former, and so $v$ must be at least a $(K+1)$th nearest neighbor of $u$. But then $r<\min_{v\in V}\left\lVert v-\operatorname{NN}_{k+1}(v)\right\rVert<\left\lVert u-v\right\rVert$, so we have a contradiction.
The proof of reverse direction nearly follows the definition of an effective connection radius. Suppose $r\geq\min_{v\in V}||v-\operatorname{NN}_{k+1}(v)||$. Let $v^{*}$ be the minimizing vertex in the expression above. Then $(v^{*},\operatorname{NN}_{k+1}(v))$ is a non-edge and $\left\lVert v^{*}-\operatorname{NN}_{k+1}(v^{*})\right\rVert\leq r$. $\blacksquare$
Our proof of Prop. 1 is loosely inspired by the proof for Lemma 58 presented by Karaman and Frazzoli [16].
@@
We prove our bound by controlling the number of samples that land within a ball of radius $r$ of another sample using a Chernoff bound.
@@
$\operatorname{\mathrm{P}}\left(\forall v\in V,\sum_{w\in V\backslash\{v\}}I_{v,w}(r)<K+1\right)$ (12)
@@
The second line takes the union bound over all vertices. The third line applies the bound $\operatorname{\mathrm{P}}(B_{r}^{d}(v)\cap X)\leq\operatorname{\mathrm{P}}(B_{r}^{d}(v))$. To obtain a tail estimate on the binomial random variable above, we use a Chernoff bound stated by Boucheron et al. [3]:
@@
where $D_{KL}$ is the KL-divergence between two Bernoulli variables, and $p=\operatorname{\mathrm{P}}(B_{r}^{d})$ and $t=K/(n-1)-p$ (our bound on $r$ will avoid a degenerate $t\leq 0$). Using the well-known inequality $D_{KL}(p_{1}||p_{2})\geq(p_{1}-p_{2})^{2}/2p_{1}$ when $p_{1}>p_{2}$:
@@
What we have shown is that if we choose a ball radius $r$ as expressed in Equation 13, then each ball of radius $r$ centered at each vertex in the PRM contains less than $K+1$ other vertices with probability at least $1-\gamma$. By the construction of the event in Eq. 12, we know that:
@@
The conservative $\log N$ factor inside the radical may be removed with a tighter analysis that does not rely on a union bound over all vertices.
A.4 Numerical Algorithm Correctness Proof and Additional Results
@@ A.4.1 Analysis of Correctness of Algorithm 2
We begin by stating the following theorem that formalizes the correctness of Algorithm 2 and then sketch out the proof:
@@
We observe from the form of the expression (lower order polynomial and exponential on the left term, higher order polynomial and linear factor on the right term) and conclude two possibilities:
  1. 1.
    $f^{\prime}(m)<0$ for all $m\geq 1$ ($f$ is monotonically decreasing), or
  2. 2.
    $f^{\prime}(m)\geq 0$ for some $k>m$, and then $f(m)<0$ for all $m>k$ ($f$ is monotonically increasing until $k$, and then starts monotonically decreasing).
If we are in the first case, we write $k=0$ for the convenience of the rest of the analysis. In either case, the doubling search of the numerical algorithm will double $N_{u}$ until $N_{u}>k$ and $f(N_{u})\leq\gamma$. Thus, we observe that the binary search will preserve two invariants:
  1. 1.
    $N_{l}\leq m^{*}\leq N_{u}$. The left inequality is true because the search increases $N_{l}$ if either $f(\left\lfloor(N_{l}+N_{u})/2\right\rfloor)>\gamma$ or $N_{l}\leq k$ (by the check that verifies $f$ is decreasing via the left inequality on Line 6). The right inequality is true via the right inequality on Line 6.
  2. 2.
    The intervals close by half each time.
Thus, the algorithm will return the correct answer since we are searching over a discrete set of integers while the invariants hold.
A.4.2 Numerical Algorithm for Proposition 1 and Tightness Experiments